PROBLEM I. Suppose that, in a market of a certain product, there is a single dominant firm with a cost function C(QcQ, where c > 0 is a constant, and the competitive fringe with a supply function Qf(p) = p-120 The market demand function is given by QM(p) = 600-3p. Q1. When c= 100, the dominant firms profit-Inaximizing quantity is (a) 100 (b) 144 (c) 180 (d) 160 (e) 120 Q2. When c- 100, the equilibrium market price of the product is (a) 120 (b) 140 (c) 180 (d) 210 (e) 160 Q3. Whenc 100, what is the market share of the dominant firm? (Choose the closest (a) 89% (b) 92% (c) 80% (d) 95% (e) 97%
Q4. Under what condition of c does the competitive fringe produce nothing at the equilibrium? (a) c^ 40. (b) c 90. (c) c2 100. (d) 100 < c < 40. (e) c 60.
Answer #1

1) The profit maximizing output of the dominant firm is given by the equation

Q= QM -Qf = 600-3p -(p-120)

                 = 720-4p

or, p= 180-Q/4 .................(i)

Therefore, total revenue of the dominant firm is,

pQ= 180Q-Q2 /4

and marginal revenue of the dominant firm is

d(pQ)/dQ= 180-Q/2

Also cost function of the dominant firm is

   TC= cQ=100Q (given c=100)

Therefore marginal cost of the dominant firm is

     MC= 100

The profit maximizing condition of a firm is

      MR=MC

or, 180-Q/2=100 or, Q =160

Therefore, the correct answer is (d)

2) The profit maximizing price of the dominant firm

Using eqtn (i) of part 1 we get,

p =180-Q/4 =180-160/4=140 which is also equilibrium market price of the product.

Therefore, the correct answer is b)

3) The market demand of the product at equilibrium market price of the product is

QM = 600-3p =600-3*140=180

Therefore, market share of the dominant firm for c=100 is

Market share = Profit maximizing output/market demand of the product*100= 160/180*100= 88.88%= 89% (Rounded to next whole integer)

Therefore, the correct answer is a)

4) From part 1) the profit maximizing condition of a firm is

    180-Q/2=c

or, 180-0/2=c (As output of the dominant firm =0 )

or, c=180

Therefore, the correct answer is c)